# Organizing Unstructured Image Collections using Natural Language

### v4

@@ 4.2 Semantic Grouper
We explore three Grouper designs that partition the dataset $\mathcal{D}$ into semantic clusters by conditioning the partition on each criterion $R_{l}\in\mathcal{R}$. The end result is a partition (or substructure) $\mathcal{O}_{l}=\{\mathcal{C}_{k}^{l}=\bigl{(}s_{k}^{l},\mathcal{D}_{k}^{l}\bigr{)}\}_{k=1}^{K_{l}}$, where an image is not only assigned to a cluster, but also has an associated semantic name (e.g., “Climbing gym” instead of a cluster index), making it human interpretable. Each design differs in how we condition the predictor to incorporate the grouping criterion.
@@
Since data can be clustered at varying granularity, we design the Grouper to cluster images at three granular levels: coarse, middle, and fine. As an example, under the Cuisine criterion, the Grouper can cluster $\mathcal{D}$ at a coarse continental level (“European”, “Asian”), a middle regional level (“Mediterranean”, “Southeast Asian”), or a fine national level (“Italian”, “Thai”).
Image-based Grouper. We task an LLM to generate a question $q_{l}$ conditioned on $R_{l}$. For e.g., for $R_{l}=\leavevmode\nobreak$Mood, the $q_{l}=$ “What mood does this image convey? Return an abstract, common, and specific category.” We then query a visual question answering (VQA) model (Li et al., 2023a) model to output a coarse-, mid-, and fine-grained label for every image as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{VQA}(\mathbf{x}_{n},q_{l})$. Subsequently, aggregating these labels over $\mathcal{D}$ at each level yields multi-granularity substructures.
Tag-based Grouper: We first prompt a LLM to generate a list of middle-grained tags (e.g., “Recreational facility”) conditioned on a given criterion $R_{l}$. Following Liu et al. (2024d), we then query the LLM to generate super- and sub-categories (e.g., “Indoor” and “Climbing gym”) for each middle-grained tag, obtaining coarse- and fine-grained tags. Given the vocabulary of tags, we use CLIP (Radford et al., 2021) to tag an image with the most likely label, one at each granular level.
@@
Caption-based Grouper (Main): We condition a MLLM with a given criterion $R_{l}$ to generate a set of captions for each image as $e_{n}^{l}=\text{MLLM}(\mathbf{x}_{n},R_{l})$. Next, we use a LLM in a three-step process to group images to clusters at multiple granularity levels: i) Initial Naming: The LLM assigns a class name to each caption as $s_{n}^{l}=\text{LLM}(e_{n}^{l},R_{l})$, producing an initial set of names $\mathcal{S}_{\text{init}}^{l}$; ii) Multi-granularity Cluster Refinement: The LLM refines $\mathcal{S}_{\text{init}}^{l}$ into three structured granularity levels as $\left(\mathcal{S}_{\text{coarse}}^{l},\mathcal{S}_{\text{mid}}^{l},\mathcal{S}_{\text{fine}}^{l}\right)=\text{LLM}(\mathcal{S}_{\text{init}}^{l},R_{l})$, which serve as candidate cluster names; iii) Final Assignment: LLM assigns each image $\mathbf{x}_{n}$ to a cluster by linking its criterion-specific caption to the structured class names each granularity level as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{LLM}(e_{n}^{l},\mathcal{S}_{\text{coarse}},\mathcal{S}_{\text{mid}},\mathcal{S}_{\text{fine}})$.
@@ 5 Experiments
Evaluation metric for criteria discovery. We assess criterion discovery performance of the proposers using the True Positive Rate (TPR) (Csurka et al., 2024), defined as $\text{TPR}=\frac{\mid\mathcal{R}\cap\mathcal{Y}\mid}{|\mathcal{Y}|}$, i.e., the fraction of ground-truth criteria $\mathcal{Y}$ recovered by the predicted set $\mathcal{R}$. Because the space of plausible criteria is inherently open-ended and subjective, False Positives are hard to define, so we use TPR as our primary metric: higher scores indicate more complete coverage of the reference annotated criteria.
Evaluation metrics for substructure uncovering. We measure how well each criterion-specific substructure predicted by the Grouper matches the ground truth along two complementary axes: i) Semantic consistency: We quantify the semantic agreement between predicted cluster names and ground-truth labels by computing their cosine similarity with Sentence-BERT (Reimers and Gurevych, 2019) for each image; the dataset-level mean is reported as Semantic Accuracy (SAcc) (Conti et al., 2023; Liu et al., 2024e). ii) Structural consistency: We quantify the structural agreement between predicted and ground-truth clusters with clustering accuracy (CAcc) (Vaze et al., 2022), which computes Hungarian-matched accuracy between the two partitions (Kuhn, 1955). Because ground-truth granularity is unknown during OpenSMC evaluation, we select the predicted substructure with the highest CAcc for evaluation. Unlike TCMC methods, which use the true cluster count during both training and evaluation, our protocol offers a fair and practical benchmark for OpenSMC.
Figure 4: Comprehensiveness Comparison of Criteria Proposers: TPR performance of each proposer is evaluated against Basic and Hard ground-truth criteria, and visualized using a Progress Bar Chart. Each block represents one ground-truth criterion, with Colored blocks indicating successfully discovered criteria and Gray blocks representing undiscovered criteria.
@@ 5.1 Study of the Criteria Proposer
Expanding ground-truth criteria for rigorous evaluation. The annotated basic criteria may not capture all valid groupings in complex datasets such as COCO-4c. We engaged human annotators to expand additional criteria for the six benchmarks in § 3, obtaining {10, 4, 11, 7, 17, 11} criteria for {Fruit-2c, Card-2c, Action-3c, Clevr-4c, COCO-4c, Food-4c}, respectively. We call the original per-image annotated criteria (Figure 2) the Basic ground truth and the enriched set the Hard ground truth. Full annotation details appear in App. C.2.
Which Proposer performs best? Figure 4 compares the TPR performance for all Proposer variants. The Caption-based Proposer recovers the most comprehensive set of criteria, ranking closest to the human-annotated reference set on every benchmark and on both Basic and Hard splits; in the latter it outperforms the runner-up Tag-based Proposer by $+32.2\%$ TPR. Captions offer richer, nuanced cues that help the LLM discover grouping criteria more comprehensively, whereas tags provide little context and therefore struggle on complex datasets (e.g., COCO-4c). The Image-based Proposer lags behind, as reasoning over small image grids omits fine visual cues and obscures overarching themes.
Impact of image quantity on criterion discovery. Figure 5 shows the TPR performance of the Caption-based Proposer as the number of input images for criteria discovery varies. Interestingly, for object-centric datasets (Card-2c, Clevr-4c), only a handful of samples – even a single image – recover most criteria, thanks to their uniform structure, i.e., seeing one playing card is enough to suggest criteria like Suit and Rank in Card-2c. In contrast, performance drops sharply on the more diverse and realistic datasets (COCO-4c, Food-4c, Action-3c) when fewer images are provided, because inferring their richer and intricate thematic criteria requires wider visual coverage. Since $\mathcal{X}$-Cluster assumes no dataset priors, we use all images by default.
@@
Figure 6: Comparison of Semantic Groupers. We report CAcc, SAcc, and their Harmonic Mean (HM) for different Semantic Groupers (Refer to caption) on the Basic criteria across six benchmarks. CLIP zero-shot classification (Refer to caption) serves as a pseudo upper bound, while KMeans (Refer to caption) with strong visual features is used as a CAcc baseline. Best performers are marked with Refer to caption. See App. G.2 for expanded numerical results and App. H for clustering visualizations.
Table 2: Comparison with TCMC methods. For each benchmark, we report the average CAcc (%) and SAcc (%) across all criteria. We provide CLIP L/14 zero-shot performance as the pseudo upper-bound reference (UB). Note: -marked methods used the ground-truth criteria and the number of clusters ($K_{l}$) as prior input. MMaP and MSub do not build semantic clusters. See expanded results in App. G.3.
@@
MMaP  (Yao et al., 2024) 33.9 - 43.8 - 62.8 - 60.6 - 36.9 - 51.0 - 48.2 -
MSub  (Yao et al., 2025) 36.0 - 47.3 - 72.2 - 64.3 - 39.6 - 54.4 - 52.3 -
@@
5.2 Study of the Semantic Grouper
Which Semantic Grouper performs best? In Figure 6, we score each Grouper with CAcc (structural agreement) and SAcc (semantic agreement) evaluated on the ground-truth per-criterion annotation, ranking models by their harmonic mean Harmonic Mean (HM) of CAcc and SAcc. To gauge headroom we include a pseudo upper bound: CLIP ViT-L/14 in a zero-shot regime where the true criterion, cluster names and count are all supplied. We also report $k$-means baselines on features from CLIP-L/14, DINOv1-B/16 (Li et al., 2023b), and DINOv2-G/14 (Oquab et al., 2024), evaluated with CAcc and given the oracle $k$.
Figure 6 shows the Caption-based Grouper performs best. It ranks first on 12/19 criteria and averages $59.9\%$ CAcc, virtually matching the upper bound’s $58.1\%$. Its mean SAcc is $60.5\%$, again best among learned methods but below the upper-bound $74.2\%$, which benefits from exact ground-truth names. This gap is expected given lexical variability, e.g., “Joyful,” “Happy,” and “Cheerful” all describe the Mood criterion but lack full semantic equivalence. The BLIP-2 Image-based Grouper ranks second. Its criterion-specific questions help, yet per-image answers add noise. The Tag-based Grouper trails across all benchmarks, reflecting the weak alignment between generated tags and dataset concepts.
Effect of multi-granularity refinement. We ablate the Caption-based Grouper with three cluster naming schemes to study the effect of our multi-granularity design: i) Initial Names, where the initially LLM-assigned names are used as the final output; ii) Flat Refinement, where the LLM refines initial names into a single-level list with uniform granularity; and our iii) Multi-Granularity Refinement strategy. As Figure 7 indicates, either refinement method markedly boosts accuracy over raw initial names, confirming the need for consistent granularity. Furthermore, our multi-granularity strategy outperforms the flat one by offering several detail levels, providing greater flexibility in aligning with user-preferred grouping granularity.
Figure 7: Study of multi-granularity refinement.
@@ 5.3 Comparison with TCMC Methods
Additional ablations and studies of $\bm{\mathcal{X}}$-Cluster. In the technical appendix, we provide extensive ablations and follow-up experiments: i) Cluster visualizations and qualitative analysis (App. H); ii) Failure case analysis (App. I); iii) Handling of hallucinated criteria (App. K); iv) Mitigating model biases (App. L); v) Computational cost analysis (App. M); vi) Sensitivity analysis of different MLLMs and LLMs (App. N); vii) Further investigation on multi-granularity clustering (App. J); and (viii) Fine-grained criterion handling (App. O).
Figure 8: Bias Discovery in T2I-Generated Images. Bias intensity, dominant clusters, and example images are shown for few occupations. Full results for all studied occupations are provided in App. P.1.
@@ 6 Applications
In this section, we apply the best caption-based $\mathcal{X}$-Cluster on three applications to showcase its ability to surface novel, human-interpretable insights.
@@
Discovering Novel Bias in Text-to-Image (T2I) Diffusion Models: Do T2I models encode biases beyond the well-studied gender and race stereotypes? (Naik and Nushi, 2023) To investigate this, we selected nine occupations (e.g., Nurse, CEO) from prior studies (Bianchi et al., 2023; Bolukbasi et al., 2016) and generated 100 images per occupation via the prompt “A portrait photo of a <OCCUPATION>” using DALL-E3 (Betker et al., 2023) and SDXL (Podell et al., 2024), resulting in 1.8k images. Applying $\mathcal{X}$-Cluster, we automatically identified 10 grouping criteria (bias dimensions) and their distributions for each occupation. To quantify bias, we measured the normalized entropy of each dimension’s distribution (D’Incà et al., 2024) as bias intensity and marked the dominant (largest) cluster as the potential bias direction. We conducted a user study with 54 participants to validate our results: $\mathcal{X}$-Cluster’s predicted bias intensity closely matched ratings from the interviewees with an Absolute Mean Error of 0.1396 on a 0–1 scale, and bias directions agree 72.3% of the time with the interviewees. See App. P.1 for user study details.
@@
Findings: As shown in Figure 8, our method identifies both well-known and previously unnoticed novel biases without relying on predefined categories. For SDXL, roles such as Nurse, Firefighter and Basketball Player show pronounced gender and racial skews that exceed labor-force statistics (U.S. Bureau of Labor Statistics, 2021), whereas DALL-E3’s guardrails (OpenAI, 2022b) mitigate many of these effects. More notably, Figure 8(d–f) highlights previously unnoticed bias dimensions. For example, SDXL strongly associates CEOs with “Grey” hair, while DALL·E3 favors “Dark” hair. Additionally, DALL·E3 shows stronger biases in Hair style and Grooming for occupations like Nurse (Figure 8(e)) and Teacher (Figure 8(f)). These findings suggest that while industrial T2I models with guardrails may have mitigated well-known biases, they often overlook emerging or less-discussed ones – highlighting the need for broader audits, where $\mathcal{X}$-Cluster can serve as a valuable tool. See further findings in App. P.1.

### v5

@@ 4.2 Semantic Grouper
Each discovered criterion $R_{l}\in\mathcal{R}$ serves as a thematic indicator for a distinct semantic substructure $\mathcal{O}_{l}$ within the image set $\mathcal{D}$. To uncover these substructures, as shown in Figure 3(right), the Grouper takes $\mathcal{D}$ and each criterion $R_{l}$ as inputs, discovers cluster names $\{s_{k}^{l}\}_{k=1}^{K_{l}}$, and groups images $\mathcal{D}_{k}^{l}$ to their corresponding clusters. As a result, the interpretable substructure $\mathcal{O}_{l}=\{\mathcal{C}_{k}^{l}=\bigl(s_{k}^{l},\mathcal{D}_{k}^{l}\bigr)\}_{k=1}^{K_{l}}$ emerges for each $R_{l}$. The core design of the Grouper focuses on aligning semantic substructure discovery with the given partitioning criterion. Like the Proposer, we explore three distinct approaches for the Grouper.
@@
Furthermore, as clusters under a given criterion can be formed at varying semantic granularities based on user preferences, we have designed our Grouper to clusters $\mathcal{D}$ at three levels: coarse, middle, and fine-grained. This allows $\mathcal{X}$-Cluster to provide insights at different granularities. For example, under the Cuisine criterion, $\mathcal{X}$-Cluster can organize images at a coarse continental level (e.g., “European” or “Asian”), a middle regional level (e.g., “Mediterranean” or “Southeast Asian”), or a fine national level (e.g., “Italian” or “Thai”). See Supp. F.2 for design details.
dlgCaption-based Grouper (main):
@@
Given a target criterion $R_{l}$, we prompt the MLLM to generate criterion-specific captions that focus exclusively on the visual content relevant to $R_{l}$ for each image, as $e_{n}^{l}=\text{MLLM}(\mathbf{x}_{n},R_{l})$. Next, we design a Multi-granularity Group Assignment (MGA) module that uses the LLM to group images into clusters across multiple semantic granularity levels through a three-step process: i) Initial Naming: The LLM assigns a provisional class name to each caption as $s_{n}^{l}=\text{LLM}(e_{n}^{l},R_{l})$, producing an initial set of names $\mathcal{S}_{\text{init}}^{l}$; ii) Multi-granularity Cluster Refinement: The LLM refines $\mathcal{S}_{\text{init}}^{l}$ into three structured granularity levels: $\left(\mathcal{S}_{\text{coarse}}^{l},\mathcal{S}_{\text{mid}}^{l},\mathcal{S}_{\text{fine}}^{l}\right)=\text{LLM}(\mathcal{S}_{\text{init}}^{l},R_{l})$, which serve as candidate cluster names; iii) Final Assignment: LLM assigns each image $\mathbf{x}_{n}$ to a cluster by linking its criterion-specific caption to the structured class names at different granularity levels as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{LLM}(e_{n}^{l},\mathcal{S}_{\text{coarse}}^{l},\mathcal{S}_{\text{mid}}^{l},\mathcal{S}_{\text{fine}}^{l})$. By aggregating these cluster assignments across $\mathcal{D}$ at different levels, we derive multi-granularity semantic substructures. As we will show in § 5.3, the Caption-based Grouper outperforms other alternatives, making it our main method.
dlgTag-based Grouper (alternative):
Given a target criterion $R_{l}$, we prompt the LLM to generate a set of common categories (e.g., “Commercial Space”) related to the criterion $\mathcal{S}_{\text{mid}}^{l}=\text{LLM}(R_{l})$ as the mid-grained tags. Following liu2024shine, we further query the LLM to infer potential super- and sub-categories (e.g., “Indoor” and “Restaurant”) for each mid-grained tag, thereby obtaining the corresponding coarse- and fine-grained tag sets, $\mathcal{S}_{\text{coarse}}^{l}$ and $\mathcal{S}_{\text{fine}}^{l}$. Finally, we employ an open-vocabulary image tagger [radford2021learning] to assign the most relevant tag at each granularity level to each image, $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{Tagger}(\mathbf{x}_{n},\mathcal{S}_{\text{coarse}}^{l},\mathcal{S}_{\text{mid}}^{l},\mathcal{S}_{\text{fine}}^{l})$, yielding multi-granularity substructures after aggregation.
dlgImage-based Grouper (alternative):
Given a target criterion $R_{l}$, we first prompt a LLM to generate a question $q_{l}$ tailored to $R_{l}$, For e.g., for the criterion Mood the generated question is: “What mood is conveyed by this image? Answer with an abstract, common, and specific category name, respectively”. We then use $q_{l}$ to guide a visual question answering (VQA) model [li2023blip] in directly inferring semantic cluster names and assignments for each image at different granularity levels as $\left(s_{\text{coarse}}^{l},s_{\text{mid}}^{l},s_{\text{fine}}^{l}\right)=\text{VQA}(\mathbf{x}_{n},q_{l})$.
@@ 5.1 Experimental Protocol
Evaluation Metric for Criteria Discovery: We use True Positive Rate (TPR) [csurka2024could] to evaluate the criteria discovery performance of different proposers. Specifically, we compute TPR as $\text{TPR}=\frac{\mid\mathcal{R}\cap\mathcal{Y}\mid}{|\mathcal{Y}|}$, measuring to what extent the predicted set covers the ground-truth criteria $\mathcal{Y}$. It is important to note that the number of grouping criteria is subjective and can be as extensive as one’s preferences allow (open-ended), making False Positives hard to define. Thus, we use TPR as the primary metric. A higher TPR means better coverage of predicted criteria compared to the ground truth.
Evaluation Metrics for Substructure Uncovering: To assess each criterion-specific substructure uncovered by the Grouper, we evaluate its alignment with the ground-truth substructure along two dimensions: i) Semantic Consistency: For each image, we compute the semantic similarity between its assigned cluster name and the ground-truth label under the current criterion using Sentence-BERT. The average similarity across the dataset, reported as Semantic Accuracy (SAcc) [liu2024democratizing], measures how well the predicted substructure aligns semantically with the ground truth. ii) Structural Consistency: We use clustering accuracy (CAcc) [han2021autonovel, vaze2022generalized] to measure the degree of structural match between the predicted and ground-truth substructures (clusters) using Hungarian matching algorithm [kuhn1955hungarian].
Since the granularity of ground-truth annotations is unknown during OpenSMC evaluation, we select the predicted substructure with the highest CAcc for assessment. Unlike TCMC methods [kwon2023image, yao2025customized] that rely on ground-truth cluster counts for perfect matching, our strategy provides a fair and practical evaluation for open-ended OpenSMC systems.
Figure 4: Comprehensiveness Comparison of Criteria Proposers. TPR performance of each proposer is evaluated against Basic and Hard ground-truth criteria, and visualized using a Progress Bar Chart. Each block represents one ground-truth criterion, with Colored blocks indicating successfully discovered criteria and Gray blocks representing undiscovered criteria.
@@ 5.2 Study of the Criteria Proposer
We also evaluate the performance of our design for the Proposer module. To properly assess his effectiveness, we realize that for complex datasets like COCO-4c, four ground-truth criteria may not cover all valid grouping options. Therefore, we expanded the ground-truth criteria for each of the six benchmarks in Sec. 3 using human annotators, resulting in {10, 4, 11, 7, 17, 11} distinct criteria for {Fruit-2c, Card-2c, Action-3c, Clevr-4c, COCO-4c, Food-4c}. We refer to the original per-image annotated criteria set (see Figure 2) as Basic ground truth and the expanded set as Hard during evaluation. See Supp. D.2 for annotations.
Which Criteria Proposer Performs the Best? In Figure 4, we compare different approaches for the Proposers in terms of the comprehensiveness of the discovered criteria using TPR. From Figure 4, we observe that our caption-based Proposer discovers the most comprehensive criteria, making it the closest to the human-annotated set among all methods. It consistently outperforms other variants in both the Basic and Hard sets across all six benchmarks. Its superior performance is particularly evident under the Hard criteria set, where it surpasses the second-best Tag-based Proposer by +32.2% TPR. Intuitively, the Caption-based Proposer works better because captions capture more diverse and nuanced aspects of the image set, which further guides the LLM to comprehensively discover different grouping criteria. Contrarily, the Tag-based Proposer is less effective in complex benchmarks (e.g. COCO-4c and Action-3c) since tags provide less contextual and descriptive information. Similarly, the Image-based Proposer is subpar in terms of performance since it is limited to reasoning over a small subset of images and loses visual details when combining images into a grid.
@@
Figure 6: Comparison of Semantic Groupers. We report CAcc, SAcc, and their Harmonic Mean (HM) for different Semantic Groupers (Refer to caption) on the Basic criteria across four benchmarks. CLIP zero-shot classification (Refer to caption) serves as an oracle, while KMeans (Refer to caption) with strong visual features is used as a CAcc baseline. The dlgbest performer for each criterion, determined by HM, is highlighted in green. Overall, our Caption-based Grouper performs best, ranking first in 10 of 15 evaluated criteria. See Supp. H.2 for clustering visualizations.
Impact of Image Quantity on Criteria Discovery: Figure 5 shows the TPR performance of the Caption-based Proposer across different image scales. Interestingly, in object-centric benchmarks like Card-2c and Clevr-4c, satisfactory performance is achieved with just a few images. In fact, even a single image often suffices for reasonable criteria discovery, as object-centric datasets tend to have uniform structures, i.e., seeing one playing card is enough to suggest criteria like Suit. However, this does not hold for more complex datasets like COCO-4c, Food-4c, and Action-3c, which feature diverse and realistic scenarios. Here, reducing the number of images leads to a clear drop in TPR performance, as capturing intricate and varied thematic criteria requires a larger image set. Since $\mathcal{X}$-Cluster operates without prior knowledge of the dataset, we use the entire dataset by default to ensure comprehensive discovery.
5.3 Study of the Semantic Grouper
Table 2: Comparison with TCMC methods. For each benchmark, we report the average CAcc (%) and SAcc (%) across all criteria. We provide CLIP L/14 zero-shot performance as the pseudo upper-bound reference (UB). Note: -marked methods used the ground-truth criteria and the number of clusters ($K_{l}$) as prior input. MMaP and MSub do not build semantic clusters. See expanded results in Supp. H.3.
@@ 5.3 Study of the Semantic Grouper
MMaP [yao2024multi] 33.9 - 43.8 - 62.8 - 60.6 - 36.9 - 51.0 - 48.2 -
MSub [yao2025customized] 36.0 - 47.3 - 72.2 - 64.3 - 39.6 - 54.4 - 52.3 -
@@
Which Semantic Grouper Performs the Best? In Figure 6, we evaluate different design choices for the Grouper using CAcc and SAcc for each criterion, determining the best performer based on Harmonic Mean (HM). To contextualize performance, we establish an oracle using CLIP ViT-L/14 in a zero-shot classification setup, where grouping criteria, cluster names, and the number of clusters are all known. We also use KMeans with ground-truth cluster numbers and visual features from CLIP-L/14, DINOv1-B/16 [li2022emergent], and DINOv2-G/14 [oquab2023dinov2] as CAcc baselines.
From Figure 6, we observe that the proposed Caption-based Grouper performs best, ranking first in 10 out of 15 tested criteria based on the HM across four benchmarks. It achieves an average CAcc of 59.9%, closely matching the oracle performance of 58.1%, highlighting the effectiveness of our text-driven approach. For SAcc, the Caption-based Grouper achieves an average of 60.5%, surpassing its counterparts, but falling short of the oracle 74.2% which benefits from exact ground-truth class names. This gap is expected due to the open nature of the semantic space, i.e., terms like “Joyful”, “Happy”, and “Cheerful” often describe the same Mood but lack full semantic equivalence. The BLIP-2 Image-based Grouper ranks second. Its criterion-specific questions improve labeling accuracy, but per-image predictions can introduce noise in clustering.
Figure 7: Ablation study of multi-granularity refinement.
Necessity of Multi-Granularity Cluster Refinement: To evaluate the effectiveness of multi-granularity cluster refinement design, we conduct controlled experiments using our Caption-based Grouper with three cluster naming strategies: i) Initial Names, where the initially assigned names are used as the final output; ii) Flat Refinement, where the LLM refines initial names into a single-level list with uniform granularity; and iii) Multi-Granularity Refinement, our proposed approach. As shown in Figure 7, both refinement methods significantly improve clustering accuracy compared to using noisy initial names, highlighting the importance of granularity-consistent cluster names for revealing substructures. Moreover, our multi-granularity refinement outperforms flat refinement by enabling clustering at different levels of detail, providing greater flexibility in aligning with user-preferred grouping granularity.
@@ 5.4 Comparison with TCMC Methods
Further Analysis of $\bm{\mathcal{X}}$-Cluster is provided in the supplementary material: i) Supp. I presents qualitative results; ii) Supp. J examines failure cases; iii) Supp. L explores how $\mathcal{X}$-Cluster handles invalid (hallucinated) criteria; iv) Supp. M investigates model biases; v) Supp. N analyzes computational costs; vi) Supp. O studies system sensitivity to different MLLMs and LLMs; vii) Supp. K further investigates the impact of multi-granularity clustering; and viii) Supp. P explores improvements for handling fine-grained criteria.
Figure 8: Bias Discovery in T2I-Generated Images. Bias intensity, dominant clusters, and example images are shown for few occupations.
@@ 6 Applications
We apply $\mathcal{X}$-Cluster to three applications, demonstrating its ability to generate novel, human-interpretable criteria for real-world analysis. Below, we present results for the first two applications, while additional results for the third application on confirming and mitigating gender bias in 162k CelebA [liu2015deep] images are provided in Supp. Q.3.
6.1 Discovering Biases in T2I Diffusion Models
@@
Do T2I models exhibit biases beyond the widely studied ones, such as gender and racial stereotypes? [naik2023social, nicoletti2023bias] To investigate this, we selected nine occupations (e.g., Nurse, CEO) from prior studies [bianchi2023easily, bolukbasi2016man] and generated 100 images per occupation using the prompt “A portrait photo of a <<OCCUPATION>>” with DALL-E3 [betker2023improving] and SDXL [podell2023sdxl], resulting in 1.8k images. Applying $\mathcal{X}$-Cluster, we automatically identified 10 grouping criteria (bias dimensions) and their distributions for each occupation. To quantify bias, we measured the normalized entropy of each distribution [d2024openbias] as bias intensity and identified the dominant cluster (the largest group) as the potential bias direction. We conducted a user study with 54 participants to validate our findings. $\mathcal{X}$-Cluster’s predicted bias intensity closely matched human ratings with an Absolute Mean Error of 0.1396 (0–1 scale) and aligned with human-identified bias directions 72.3% of the time. User study details are provided in Supp. Q.1.
@@
Findings: As shown in Figure 8, our method identifies both well-known and novel biases in occupational images without relying on predefined categories. For instance, Figure 8(a–c) reveals strong gender and racial imbalances in SDXL-generated images for roles like Nurse, Firefighter, and Basketball Player, exceeding official statistics [blsEmployedPersons]. In contrast, DALL·E3 exhibits improved bias mitigation, likely due to its built-in “guardrails” [dalle_guards]. More notably, Figure 8(d–f) highlights previously unrecognized bias dimensions. For example, SDXL strongly associates CEOs with “Grey” hair, while DALL·E3 favors “Dark” hair. Additionally, DALL·E3 shows stronger biases in Hair style and Grooming for occupations like Nurse (Figure 8(e)) and Teacher (Figure 8(f)). These findings suggest that while industrial T2I models with guardrails may address well-known biases, they may still overlook emerging or less-discussed ones, underscoring the need for broader bias analysis. For additional findings and experimental details, see Supp. Q.1.
6.2 Analyzing Social Media Image Popularity